Find the linear substitution that reduces the inhomogeneous equation to the homogeneous one: (8y + 10x)dx +(7x +5y)dy = 0.


1
Expert's answer
2021-11-26T00:04:00-0500
"(8y + 10x)dx +(7x +5y)dy = 0"

"(7x +5y)y' =-(8y + 10x)"

"y'=-\dfrac{8y+10x}{7x+5y}"

Use the linear substitution  "y=vx, v" is a function of "x"


"y'=v+v'x"

"v+v'x=-\dfrac{8vx+10x}{7x+5vx}"

The homogeneous

"v'x=-\dfrac{13v+17}{7+5v}"

"\dfrac{5v+7}{13v+17}=-\dfrac{dx}{x}"